Solve by the elimination method
2x + 3 y = 2
4x + 6y = 4
The solution is ????? (ordered pair)

Answer by stanbon(75887)

Solve by the elimination method
2x + 3 y = 2
4x + 6y = 4
----------------------
multiply thru the 1st equation by 2:
4x + 6y = 4
----
Subtract that from the 2nd equation and you get:
0 = 0
------------
That means the 2nd equation is simply a multiple
of the 1st. The two equations give you the same
set of points.
Each point is of the form (x, (-2/3)x + (2/3))
-----
The system is said to be dependent because one
equation is a multiple of the other.
=====================================
Cheers,
Stan H.